How much money should be put in a savings account now that earns 6.0% a year compounded weekly if you want to have dollar-sign 50,000 in 15 years?
Round your answer to two decimal places.

Answer by stanbon(75887)

How much money should be put in a savings account now that earns 6.0% a year compounded weekly if you want to have dollar-sign 50,000 in 15 years?
=====================
A(t) = P(1+(r/n))^(nt)
-----
50,000 = P(1+(0.06/52))^(52*15)
------
50,000 = P(2.4583)
----
Ans: P = 50,000/2.4583 = $20,339.03
-------------------
Cheers,
Stan H.
-------------------